# Runtime Stealthy Perception Attacks against DNN-based Adaptive Cruise Control Systems

### v3

@@ Appendix B Malware Installation
Fig. 19 shows an example set of steps taken for establishing remote access and downloading a malicious code repository on a vehicle running OpenPilot, the open-source production ACC system from Comma.ai [29]. To change live camera image frames at runtime, we use a technique called monkey patching in Python, which is used for dynamically modifying or extending the behavior of an existing code at runtime or hooking a function without changing the source code[42]. For example, as shown in Fig. 19, the send() system library that is called by the camera call back function for sending the received camera frames to the DNN model can be wrapped by a malicious version adv_send() that implements the attack.
Figure 19: An example of malware installation (Top) and malicious code execution using monkey patching technique (Bottom).
@@ Appendix C AEBS Evaluation
For a realistic implementation of AEBS in our simulation platform, we study the AEBS design of typical OpenPilot-supported car models [9]. Our AEBS design relies on the fusion of camera and radar data, as mentioned in Appendix A, aligning with AEBS design of most current OpenPilot-supported car models, as shown in Table IX.
TABLE IX: AEBS design in OpenPilot-supported car models.
Car Model [9] Is AEBS using Radar/Camera/Both
@@
Following the testing protocol specified in [14], we employ two categories and five driving scenarios (see Table X) to assess our AEBS functionality (Section 6). Each scenario is repeated 100 times to ensure reliable outcomes. Experimental results show that in all five testing scenarios, both FCW and AEB alerts are activated, effectively preventing all hazards or collisions. On average, it takes approximately 1.68 seconds for AEB to stop the Ego vehicle completely.
TABLE X: Driving scenarios to test the AEBS with different initial distances ($Init\_dist$) between the Ego vehicle and the lead vehicle.
@@ Appendix D Stealthiness User Study
We conduct a user study [2] to further evaluate the advantages of the stealthiness design of our attack. Before recruiting participants, we secured Institutional Review Board (IRB) approval. Our study explicitly avoided collecting any personally identifying information, targeting sensitive populations, or introducing any risks to the participants.
Our study included 30 participants who were asked to sit on the driver’s side of an autonomous vehicle, parked in a parking lot, equipped with OpenPilot ADAS (see Section II-C). Each participant went through different trials of pre-recorded videos displayed on the ADAS monitor and answered a series of questions after each trial using a Qualtrics survey [8]. All participants had driving experience and 40% of them had autonomous driving experience.
@@
Driving Preferences. We first ask participants to envision themselves driving this autonomous vehicle with the ADAS monitor displaying pre-recorded image frames. We inquire about how often they would look at the ADAS monitor while driving and whether alterations in the monitor’s position and size influence their preference. User study results in Fig. 20 show that 99% of the participants express a preference for looking at the ADAS monitor during their driving experience, with 33% specifying they would do so for the majority of the driving duration. Moreover, 60% of the participants indicate a preference for a larger monitor size or a more prominent position.
@@
Figure 20: Results of participants’ preference of looking at the ADAS monitor during driving and whether they would look more often at the monitor with a larger size or in a noticeable position.
Stealthiness. We create five video sets by introducing adversarial patches into a pre-recorded highway scenario using CA-Random, CA-APGD, and CA-Opt methods with three stealthiness levels ($\lambda=10^{-4}$, $\lambda=10^{-3}$, $\lambda=10^{-2}$, as detailed in Section V-C1). We present these videos on the ADAS monitor and ask participants whether they notice any abnormal scenarios that prompt them to assume control of the vehicle to avoid potential risk or danger. For a detailed examination, we extract an image frame from each video at the same frame index and zoom in to reveal more intricate details, followed by posing identical questions to the participants.
Results of the user study are illustrated in Fig. 21. It is evident that patches generated by the CA-Random attack are conspicuous to the majority of participants (>75%), whether observed in images or videos. In comparison, patches generated by the CA-APGD exhibit lower visibility than those produced by the CA-Random attacks. In CA-Opt attacks, the takeover rate diminishes with a rise in stealthiness level or $\lambda$ value. Specifically, when $\lambda$ is set at $10^{-2}$ and $10^{-3}$, the takeover rates are below 20% for patch images and are 0% for patch videos. These findings suggest that adversarial patches at $\lambda=10^{-2}$ and $\lambda=10^{-3}$ exhibit nearly imperceptible characteristics to human drivers, particularly in image frames when not zoomed in.
Figure 21: Results of stealthiness of each attack method.
Physical Attack. We also investigate the stealthiness of physical adversarial patches as perceived by human eyes. Participants are shown an image of an adversarial patch generated through a physical attack method introduced in a prior work [39]. They are then asked identical questions. Our findings reveal a takeover rate of 80%, highlighting the inadequacy of physical patches in achieving stealthiness and evading human detection.
@@ Appendix E Evaluation of Fake Video Attacks
Fake Video Attack. We rerun the simulations for each scenario introduced in Section V-A and replace the real-time camera frames with the selected fake video when the Ego vehicle approaches a similar position indicated by the fake video. Experimental results show that this attack causes hazards in 100% simulations and in 72.6% of simulations the Ego vehicle drives to the neighbor lane without any collisions. This is because lane lines in the fake video do not exactly overlap with the ones in the actual video.
@@
Due to such differences, fake video attacks can be easily detected by existing methods that monitor the differences between two consecutive image frames. An example of image frame changes during a perfect fake video attack is shown in Fig. 22. Even if the fake videos are recorded using the same camera on the Ego vehicle driving in the same lane and weather conditions with only one lead vehicle (resembling replay attacks), an alert human driver can still notice the changes in the lead vehicle’s position and size. An example of the RMSE and UIQ [86] between two consecutive image frames is shown in Fig. 23. We see that the similarity between the first frame of the fake video and the last frame of the benign video is much lower than that between other consecutive frames.
@@
Figure 22: An example of two consecutive images before (Left) and after (Right) fake video attack.
@@
Figure 23: Similarity of two consecutive image frames with CA-Opt attack and fake video attack (starting at 69th frame) measured in RMSE and universal image quality index (UIQ) [86].
@@
Fake Video Attack with Safety Interventions. To further evaluate the performance of fake video attacks with safety mechanisms, we rerun the experiments by launching the proposed driver intervention 2.5 seconds (average reaction time) after the attack. Experimental results show that all the attacks are successfully prevented. Therefore, we do not further test the fake video attack while enabling AEBS and constraint checking (see Section IV).
Appendix F Testing on a Real-world Dataset.
We perform an evaluation using the comma2k19 dataset [77], a publicly available dataset with over 33 hours of California’s 280 highway commute. The dataset comprises 2019 segments, each lasting one minute, covering a 20km highway section, collected using OpenPilot hardware.
@@
From this dataset, we choose 200 videos with a clear view of the lead vehicle and a relative distance of less than 100 meters. These videos are fed into the DNN model to record predictions for relative distance (considered as ground truth). We then introduce adversarial patches, generated by different attack methods, into the videos. The manipulated videos are fed into the DNN model, and predictions for relative distance are compared with those from videos without any attacks, using metrics like average and standard deviations in the predicted longitudinal distance.
Table XI compares the CA-Opt attack with CA-Random for different distances between the Ego and lead vehicles. CA-Random has an average deviation of 0.15m, which does not significantly impact ACC system outputs or cause hazards as the ACC system typically keeps a following distance larger than 4 meters in the absence of attacks. In contrast, when the Ego vehicle is close to the lead vehicle (less than 20m), CA-Opt achieves the highest deviation of 18.65m, showcasing the effectiveness of the proposed objective function (Section III-C2) in generating optimal perturbations with substantial impact.
@@
TABLE XI: Performance of CA-Opt attack vs. CA-Random in deviating DNN-based lead vehicle position predictions using comma2k19 dataset.

### v4

@@ Appendix B Malware Installation
Fig. 17 shows an example set of steps taken for establishing remote access and downloading a malicious code repository on a vehicle running OpenPilot, the open-source production ACC system from Comma.ai (commaai-openpilot, ). To change live camera image frames at runtime, we use a technique called monkey patching in Python, which is used for dynamically modifying or extending the behavior of an existing code at runtime or hooking a function without changing the source code(hunt2019monkey, ). For example, as shown in Fig. 17, the send() system library called by the camera call back function for sending the received camera frames to the DNN model can be wrapped by a malicious version adv_send() that implements the attack.
Figure 17. An example of malware installation (Top) and malicious code execution using monkey patching technique (Bottom).
@@ Appendix C AEBS Evaluation
For a realistic implementation of AEBS in our simulation platform, we study the AEBS design of typical OpenPilot-supported car models (Carmodels-OpenPilot, ). Our AEBS design relies on the fusion of camera and radar data, as mentioned in Appendix A, aligning with AEBS design of most current OpenPilot-supported car models, as shown in Table 9.
Table 9. AEBS design in OpenPilot-supported car models.
Car Model (Carmodels-OpenPilot, ) Is AEBS using Radar/Camera/Both
@@
Following the testing protocol specified in (http://data.europa.eu/eli/reg/2020/1597/oj, ), we employ two categories and five driving scenarios (see Table 10) to assess our AEBS functionality (Section 6). Each scenario is repeated 100 times to ensure reliable outcomes. Experimental results show that in all five testing scenarios, both FCW and AEB alerts are activated, effectively preventing all hazards or collisions. On average, it takes about 1.68 seconds for AEB to stop the Ego vehicle completely.
Table 10. Driving scenarios to test the AEBS with different initial distances ($Init\_dist$) between the Ego vehicle and the lead vehicle.
@@ Appendix D Stealthiness User Study
We conduct a user study (adassurvey, ) to further evaluate the advantages of the stealthiness design of our attack. Before recruiting participants, we secured Institutional Review Board (IRB) approval. Our study explicitly avoided collecting any personally identifying information, targeting sensitive populations, or introducing any risks to the participants. Our study included 30 participants who were asked to sit on the driver’s side of an autonomous vehicle, parked in a parking lot, equipped with OpenPilot ADAS (see Section 2.3). Each participant went through different trials of pre-recorded videos displayed on the ADAS monitor and answered a series of questions after each trial using a Qualtrics survey (qualtrics, ). All participants had driving experience, and 40% of them had AV driving experience.
@@
Driving Preferences. We first ask participants to envision themselves driving this autonomous vehicle with the ADAS monitor displaying pre-recorded image frames. We inquire about how often they would look at the ADAS monitor while driving and whether alterations in the monitor’s position and size influence their preference. User study results in Fig. 18 show that 99% of the participants prefer looking at the ADAS monitor during their driving experience, with 33% specifying they would do so for the majority of the driving duration. Moreover, 60% of the participants indicate a preference for a larger monitor size or a more prominent position.
@@
Figure 18. Results of participants’ preference of looking at the ADAS monitor during driving and whether they would look more often at the monitor with a larger size or in a noticeable position.
Stealthiness. We create five video sets by introducing adversarial patches into a pre-recorded highway scenario using CA-Random, CA-APGD, and CA-Opt methods with three stealthiness levels ($\lambda=10^{-4}$, $\lambda=10^{-3}$, $\lambda=10^{-2}$, as detailed in Section 5.3.1). We present these videos on the ADAS monitor and ask participants whether they notice any abnormal scenarios that prompt them to assume control of the vehicle to avoid potential risk or danger. For a detailed examination, we extract an image frame from each video at the same frame index and zoom in to reveal more intricate details, followed by posing identical questions to the participants.
Figure 19. Results of stealthiness of each attack method.
User study results are illustrated in Fig. 19. It is evident that patches generated by the CA-Random attack are conspicuous to the majority of participants (>75%), whether observed in images or videos. In comparison, patches generated by the CA-APGD exhibit lower visibility than those produced by the CA-Random attacks. In CA-Opt attacks, the takeover rate diminishes with a rise in stealthiness level or $\lambda$ value. Specifically, when $\lambda$ is set at $10^{-2}$ and $10^{-3}$, the takeover rates are below 20% for patch images and are 0% for patch videos. These findings suggest adversarial patches at $\lambda$ = $10^{-2}$ and $\lambda=10^{-3}$ exhibit nearly imperceptible characteristics to human drivers, particularly in image frames when not zoomed in.
Physical Attack. We also investigate the stealthiness of physical adversarial patches as perceived by human eyes. Participants are shown an image of an adversarial patch generated through a physical attack method introduced in a prior work (guo2023adversarial, ). They are then asked identical questions. Our findings reveal a takeover rate of 80%, highlighting the inadequacy of physical patches in achieving stealthiness and evading human detection.
@@ Appendix E Evaluation of Fake Video Attacks
Fake Video Attack. We rerun the simulations for each scenario introduced in Section 5.1 and replace the real-time camera frames with the selected fake video when the Ego vehicle approaches a similar position indicated by the fake video. Experimental results show that this attack causes hazards in 100% simulations and in 72.6% of simulations the Ego vehicle drives to the neighbor lane without any collisions. This is because lane lines in the fake video do not exactly overlap with the ones in the actual video.
@@
Figure 20. An example of two consecutive images before (Left) and after (Right) fake video attack.
Due to such differences, fake video attacks can be easily detected by existing methods that monitor the differences between two consecutive image frames. An example of image frame changes during a perfect fake video attack is shown in Fig. 20. Even if the fake videos are recorded using the same camera on the Ego vehicle driving in the same lane and weather conditions with only one lead vehicle (resembling replay attacks), an alert human driver can still notice the changes in the lead vehicle’s position and size. An example of the RMSE and UIQ (wang2002universal, ) between two consecutive image frames is shown in Fig. 21. We see that the similarity between the first frame of the fake video and the last frame of the benign video is much lower than that between other consecutive frames.
@@
Figure 21. Similarity of two consecutive image frames with CA-Opt attack and fake video attack (starting at 69th frame) measured in RMSE and universal image quality index (UIQ) (wang2002universal, ).
@@
Fake Video Attack with Safety Interventions. To further evaluate the performance of fake video attacks with safety mechanisms, we rerun the experiments by launching the proposed driver intervention 2.5 seconds (average reaction time) after the attack. Experimental results show that all the attacks are successfully prevented. Therefore, we do not further test the fake video attack while enabling AEBS and constraint checking (see Section 4).
Appendix F Robustness to Real-world Factors
To assess attack robustness, we vary front camera height based on standard passenger car profiles from manufacturers (urazghildiiev2007vehicle, ). We perform our experiments with four heights between 1.1-1.7 meters and three initial distances (50m, 75m, 100m). Fig. 22 illustrates the 100% success rate of our CA-Opt attack across 12 testing scenarios. The Ego vehicle initially maintains a safe following distance, deviates from it around the 2,500-3,000 control cycle or step due to the adversarial patch, and eventually collides with the lead vehicle. These results demonstrate our attack is robust to different camera positions and initial longitudinal distances and can cause safety hazards. We also test our attacks with diverse weather (rainy, sunny, or cloudy) and lighting conditions (noon or sunset). Results show that our CA-Opt attack causes longitudinal deviations of 9.8-14.3m in the predicted lead vehicle position, while maintaining a success rate of 100% under such conditions.
Figure 22. Actual relative distance trajectories under CA-Opt attack with different camera heights (H1:1.1m, H2:1.3m, H3:1.5m, H4:1.7m) and initial longitudinal distances to the lead vehicle (L1:50m, L2:75m, L3:100m). An actual relative distance of zero indicates collision.
Appendix G Runtime Overhead
Experimental results show that the time overhead introduced by the context inference component before activating the attacks is minimal (1.17 us). Following the activation of attacks, the time overhead for the object detection module is about 10.1 ms on average. Note that some production ADAS provide object detection and tracking features, so this overhead time could be potentially avoided. We also observe that the primary attribution algorithm (sundararajan2017axiomatic, ) does not add significant overhead, leveraging gradients calculated during the patch optimization process. The total time overhead is 1.52 ms.
Figure 23. Runtime overhead of each step of the attack.
Appendix H Testing on a Real-world Dataset.
@@ Appendix G Runtime Overhead
We perform an evaluation using the comma2k19 dataset (comma2k19, ), a publicly available dataset with over 33 hours of California’s 280 highway commute. The dataset comprises 2019 segments, each lasting one minute, covering a 20km highway section, collected using OpenPilot hardware. From this dataset, we choose 200 videos with a clear view of the lead vehicle and a relative distance of less than 100 meters. These videos are fed into the DNN model to record predictions for relative distance (considered as ground truth). We then introduce adversarial patches, generated by different attack methods, into the videos. The manipulated videos are fed into the DNN model, and predictions for relative distance are compared with those from videos without any attacks, using metrics like average and standard deviations in the predicted longitudinal distance.
@@
Table 11. Performance in deviating DNN-based lead vehicle position predictions using comma2k19 dataset.
@@
Table 11 compares the CA-Opt attack with CA-Random for different distances between the Ego and lead vehicles. CA-Random has an average deviation of 0.15m, which does not significantly impact ACC system outputs or cause hazards as the ACC system typically keeps a following distance larger than 4 meters in the absence of attacks. In contrast, when the Ego vehicle is close to the lead vehicle (less than 20m), CA-Opt achieves the highest deviation of 18.65m, showcasing the effectiveness of the proposed objective function (Section 3.3.2) in generating optimal perturbations with substantial impact.